1)A 9.7 L cylinder contains a sample of gas at 14.1 atm. What pressure would the gas be in a 6.5 L cylinder at this same temperature?

Report your answer in atmospheres to the correct number of significant figures.

2)A sample of gas in a 29.3 L container is initially at 15°C and 641 mmHg. If the pressure is reduced to 465 mmHg and the temperature is increased to 61°C, what is the final volume?

Add 273 to convert to Kelvin.

Report your answer in liters to three significant figures.
Answer #1

1)

Given:

Pi = 14.1 atm

Vi = 9.7 L

Vf = 6.5 L

use:

Pi*Vi = Pf*Vf

14.1 atm * 9.7 L = Pf * 6.5 L

Pf = 21.0 atm

Answer: 21.0 atm

2)

Given:

Pi = 641 mmHg

Pf = 465 mmHg

Vi = 29.3 L

Ti = 15.0 oC

= (15.0+273) K

= 288 K

Tf = 61.0 oC

= (61.0+273) K

= 334 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(641 mmHg*29.3 L)/(288.0 K) = (465 mmHg*Vf)/(334.0 K)

Vf = 46.8 L

Answer: 46.8 L